You are a financial advisor. A client of yours wants to know which of the following two options would earn more interest over 10 years: (a) Depositing $50 every month into a savings account with an annual interest rate of 6% compounded monthly, or (b) Depositing $600 every year into a savings account with an annual interest rate of 6%


1
Expert's answer
2021-05-07T09:08:45-0400

Depositing$50

"A=50\times (1-(1.005)^{-120})/.005=4,504"


Depositing$600

"A=600\times (1-(1.06)^{-10})/.06=4,416"


Depositing$50 monthly yields more than depositing$600 yearly by $ 88